OpenStudy (anonymous):

standard form of: (x)(x-3)(x-5)

OpenStudy (anonymous):

Distribute and combine like terms.

OpenStudy (anonymous):

x(x-3)(x-5) x^3-8x^2+15x x((x-8)+15) where u can get x=0, 3, &5

OpenStudy (anonymous):

(x)(x-3)(x-5)= \[x(x ^{2}-8x+15)\] \[x ^{3}-8x ^{2}+1\]

OpenStudy (anonymous):

x(x-3)(x-5) x^3-8x^2+15x x((x-8)x+15) where u can get x=0, 3, &5

OpenStudy (anonymous):

You won't get solutions.. it's not an equation.

OpenStudy (anonymous):

\[x ^{3}-8x ^{2}\+15]